# Concept Corrector: Erase concepts on the fly for text-to-image diffusion models

### v1

@@ C Reproduction Details
D Instructions for Prompt Generation
@@
E Negative Prompts
@@
Table S9: The negative prompts for the erased concepts.
Category Concept Element Negative Prompt
@@ I Time Efficiency
Tab.S16 presents the running time of the methods for generating one image. The diffusion scheduler is DDIM and the sampling step ($T$) is 50. $t_{1}$ is set to 40 and $t_{2}$ is set to 20, which follows the configurations in the experiments. The measurements are conducted on the NVIDIA A100 40GB GPU. All methods run in the same environment.
@@
Method Time in Generation Stage Total
$[T,t_{1}]$ Check-1 $[t_{1},t_{2}]$ Check-2 $[t_{2},0]$
SD v2.1 0.57 - 1.15 - 1.09 2.81
SLD-Max 0.80 - 1.62 - 1.53 3.95
SAFREE 2.91 - 5.84 - 5.55 14.30
Ours 0.60 2.50 1.68 2.46 1.24 8.48

### v2

@@ B Reproduction Details
C Instructions for Prompt Generation
@@
D Query for VLM
We design the following query in the Generation Check Mechanism for the VLM.
E Negative Concepts
@@
Table S10: The negative concepts for the erased concepts.
Category Concept Element Negative Concept
@@ I Time Efficiency
Method Generation Time Check Time Total
SD v2.1 2.81 - 2.81
SLD-MAX 3.95 - 3.95
SAFREE 14.30 - 14.30
Ours (VLM Detector) 3.52 4.96 8.48
Ours (Simple Detectors) 3.56 0.70 4.26
Table S18: The erasure and preservation performance of our method when implementing the VLM and the simple detectors for concept checking respectively.
[…]
Nude Shock Bird Couch Van. Monet CLIP
VLM 4.0 37.0 3.5 4.5 2.0 3.5 30.81
Simple 4.5 24.5 4.5 7.5 3.5 0.5 30.88
Tab.S17 presents the running time of the methods for generating one image. The experiments follow the configurations in the main paper and run on the NVIDIA A100 40GB GPU.
The results demonstrate that our method is faster than both SLD and SAFREE in generation time. The efficiency bottleneck of the whole pipeline is VLM. Considering that it only acts as a binary concept detector, it can be easily replaced with simple concept-specific detectors so that the check can be accelerated significantly while preserving the erasure performance. To demonstrate this point, we perform the extended experiments, replacing the VLM with some simple detectors. These detectors are the ones used in the evaluation. The results of their time efficiency and performance are listed in Tab.S17 and Tab.S18 respectively, showing that the check time is greatly reduced, while the erasure and preservation performance is similar or even improved further on some concepts. In order to erase unwanted concepts effortlessly, we still use the configuration of VLM as the detector to conduct the main experiments. We recommend using a lightweight concept-specific detector in practical application scenarios where a fast response speed is required.
@@
K Limitations
As a training-free approach, Concept Corrector introduces some additional operations during generation while preserving model parameters. It provides superior applicability to commercial closed-source models, yet creates implementation vulnerabilities for open-source models at the same time. Specifically, it allows malicious users with code access to potentially bypass our method through intentional code modifications.
It should be noted that it is necessary and important to focus on the commercial deployment scenario. In real-world business applications where model retraining is often impractical, our method provides essential safeguards against compliance risks. For commercial AI services, non-compliance may result in substantial financial penalties, legal repercussions, and reputation damage. In the future, we will investigate hybrid approaches combining these operations with selective parameter tuning to enhance open-source robustness.